The Ka for acetic acid, CH3COOH, is 1.8 × 10-5. A buffer, made from 0.10 M CH3COOH and 0.10 M CH3COO- has a pH of ________.

a. 4.74

b. 14.00

c. 1.00

d. 9.26

e. 7.00
Answer #1

Ka   = 1.8*10^-5

PKa   = -logKa

         = -log1.8*10^-5

          = 4.74

[CH3COOH]    = 0.1M

[CH3COO^-]    = 0.1M

PH    = PKa + log[CH3COO^-]/[CH3COOH]

          = 4.74 + log0.1/0.1

           = 4.74 + log1

           = 4.74 +0

          = 4.74 >>>>>answer
a. 4.74 >>>>answer